# From the Stochastic Embedding Sufficiency Theorem to a Superspace Diffusion Framework

### 20 Mar 2026 v1

@@ 4.6 Relation to Nelson’s Stochastic Mechanics
Nelson [6] solved the forward problem for quantum mechanics: given $\sigma=\sqrt{\hbar/m}$, derive the Schrödinger equation. The present work solves the inverse problem across nine domains: given time series data, recover $\sigma$ without assuming the physics. The forward and inverse paths converge at C4, where the pipeline recovers Nelson’s $\sigma$ to 0.026%. Nelson’s insight, that physical law admits stochastic content [24], is here extended across the $\sigma$-continuum to gravitational degrees of freedom.
@@ 4.6.1 Distinguishing the stochastic and quantum formulations
The correspondence does not, however, extend to the level of solutions and physical predictions. Three structural properties of the Fokker–Planck evolution: on which the galactic-scale predictions of Part II depend: have no counterparts in the Wheeler–DeWitt equation.
First, the Fokker–Planck equation possesses a unique normalisable steady state. For the galactic-scale equation derived in Part II (Appendix A, Theorem A.8), this steady state is the Gamma distribution $P_{\mathrm{ss}}(g)=\mathrm{Gamma}(\mathcal{E}_{r}+d_{\mathrm{eff}},\,a_{0})$, from which the equilibrium prediction $\gamma^{2}=1+d_{\mathrm{eff}}/\mathcal{E}_{r}$ follows as the ratio of the mean to the baryonic acceleration. The Wheeler–DeWitt equation $\hat{H}\Psi=0$ is a constraint equation: it has no time parameter, no spectral gap, and no mechanism for selecting a unique solution from its solution space.
Second, the Fokker–Planck evolution is dissipative. It drives any initial probability distribution toward the unique steady state on a timescale set by the spectral gap $\lambda_{1}=1/\tau_{0}$. This irreversible relaxation produces definite predictions: regardless of initial conditions, the acceleration distribution converges to the Gamma steady state. Quantum evolution preserves unitarity: pure states remain pure and the evolution is time-reversible: producing recurrences rather than relaxation.
Third, the Fokker–Planck steady state depends on the positivity of $P[g]\geq 0$ and the real-valued character of the diffusion coefficient. Under the Wick rotation, $P$ maps to a complex-valued wave functional $\Psi[g]$ that is generically oscillatory and not normalisable in the standard $L^{2}$ sense. The Gamma distribution: a positive, real-valued, normalisable function on the half-line: has no natural image under this mapping.
@@
The framework therefore proposes a specific physical hypothesis: that the gravitational degrees of freedom evolve by diffusion on superspace, governed by the Fokker–Planck equation (Axiom A2), rather than by unitary quantum evolution governed by the Wheeler–DeWitt equation. The two formulations share a generator and a classical limit (general relativity, recovered as $L\gg\ell_{P}$ in both cases) but make distinct predictions for the statistical distribution of gravitational acceleration at galactic scales. Part II subjects these predictions of the Superspace Diffusion Framework to observational data across 19 galactic probes.
@@ 4.6.2 Relationship to conventional approaches to quantum gravity
The stochastic formulation is therefore not a reformulation of any existing approach to quantum gravity, but a distinct physical theory that shares their classical limit and connects to each through specific mathematical correspondences.
@@ 4.7 Limitations
All C1–C9 data are synthetic. The pipeline has not been applied to raw experimental data in this work. The synthetic data validates the methodology, where experimental application is the subject of subsequent publications.
@@
The axioms A1–A4 are postulates, not theorems. Their justification rests on the empirical evidence of the $\sigma$-continuum, the internal consistency demonstrated in Appendices C and D, and the absence of tuneable parameters. Alternative axiomatisations may exist.
Although the inverse recovery pipeline uncovers known physics from synthetic data, the framework formalised by axioms A1–A4 generates specific quantitative predictions that can be tested against observational data. A companion paper (Part II) maintains the data-driven approach of the present work, subjecting these predictions to galactic kinematic data.
@@ 5 Conclusion
Application of the resulting pipeline to nine physical domains (classical mechanics, statistical mechanics, nuclear physics, quantum mechanics, chemical kinetics, electromagnetism, relativistic quantum mechanics, quantum harmonic oscillator dynamics, and quantum electrodynamics) recovers the known governing equations of each domain blindly, with recovery errors ranging from $0.026\%$ to $\sim\!1\%$ and no null hypothesis rejected at the 5% significance level. Fundamental physical constants emerge in both the drift and the diffusion channels without prior specification. The recovered diffusion coefficients, viewed together, constitute an empirical pattern: the $\sigma$-continuum.
Interpretation of the relativistic and quantum field theory results yields three structural consequences under the stated assumptions. First, the Dambis–Dubins–Schwarz time-change theorem combined with the Klein–Gordon density produces the relativistic suppression $\sigma=\sigma_{0}/\gamma$, with physical constants emerging in both channels. Second, the massless cancellation and gravitational self-coupling together determine $\sigma=\ell_{P}/\sqrt{t_{P}}$ for gravitational degrees of freedom, with three independent consistency arguments converging on the dimensionless prefactor $\alpha=1$. Third, a set of four canonical axioms (A1–A4) yields a framework within which physical time, the drift, the covariance operator, and the fluctuation amplitude are all fixed.
@@
is non-parametric: all coefficients are derived from the axioms.
The Superspace Diffusion Framework defined by axioms A1–A4 is falsifiable. A companion paper (Part II) maintains the data-driven approach of the present work, subjecting its predictions to observational galactic kinematic data across 19 probes spanning four physics channels.
@@ Acknowledgements
Figure 1: The $\sigma$-Continuum. Recovered diffusion coefficient $\sigma$ across all nine physical domains and the theoretical extrapolation to gravity, displayed on a logarithmic vertical axis ($\log_{10}\sigma$, schematic units). Each domain is represented by a distinct marker shape and colour, labelled with its recovered $\sigma$ expression. Deterministic domains (C1: classical mechanics, dark blue circle; C6: classical electromagnetism, light blue circle) are placed at $\log_{10}(\sigma)\approx-40$ for visualisation, labelled $\sigma=0$. Stochastic domains span intermediate scales: C2 (statistical mechanics, green square, $\sigma=\sqrt{2\gamma k_{B}T/m}$), C3 (radioactive decay, orange diamond, $\sigma^{2}=\mu$), C4 (quantum mechanics, purple pentagon, $\sigma=\sqrt{\hbar/m}$), C5 (chemical kinetics, red triangle, $\sigma\propto\Omega^{-1/2}$). The relativistic and quantum field theory domains occupy the upper region: C7 (Klein–Gordon, red–brown pentagon, $\sigma=\sqrt{\hbar/m}/\gamma$), C8 (quantum harmonic oscillator, brown diamond, $\sigma=\sqrt{\gamma\hbar\omega/m}$), C9 (QED photon, magenta circle, $\sigma=c\sqrt{\gamma}$). The dashed lilac horizontal line marks the quantum scale $\sigma\sim\sqrt{\hbar/m}$; the dashed gold horizontal line marks the Planck scale $\ell_{P}\approx 10^{-35}$ m. The gold star at the far right represents the theoretical extrapolation $\sigma=\ell_{P}$ for gravitational degrees of freedom, the inference developed in §4 and formalised in Appendices C and D. The continuum from $\sigma=0$ to $\sigma=\ell_{P}$ is an empirical pattern arising from nine independent blind recovery experiments.
@@
Figure 5: C3: Radioactive Decay ($\sigma^{2}=\mu$, Fano =1=1). (a) Poisson count series at $\mu=3000$ (blue) over 500 measurements, with the 250-realisation ensemble median (green) and 95% CI band (green shading). Inset: probability densities of the data (blue), ensemble (green), and theoretical Poisson–Gaussian approximation (red). (b) Recovered $\hat{\sigma}^{2}$ (pipeline) versus mean count $\mu$ on log–log axes (magenta circles) across seven mean count levels from $\mu\approx 100$ to $\mu\approx 10\,000$. The dashed red line is the Poisson prediction $\sigma^{2}=\mu$. The log–log slope is $1.0002$ ($R^{2}=0.9997$). (c) Observed count distribution at $\mu=3000$ (blue histogram) overlaid with the fitted Gaussian ($\mu=3001$, $\sigma=55.1$; red curve). The measured Fano factor is $1.0109$, consistent with Poisson statistics ($F=1$). (d) Autocorrelation function of the count series at 50 lags. The red dashed lines mark the 95% significance band ($\pm 0.0196$). Three of 50 lags exceed the band ($\sim$2 expected by chance), confirming temporal independence, a necessary property of Poisson counting. (e) Distribution of Fano factor $F=\hat{\sigma}^{2}/\hat{\mu}$ from 50 independent simulations at $\mu=3000$; median =1.0000=1.0000 (green line), 95% CI $[0.9612,1.0295]$, $p(H_{0}\!:F=1)=0.503$. (f) Variance $\hat{\sigma}^{2}$ versus mean $\mu$ from 50-simulation medians at seven count levels (magenta circles with 95% CI); the log–log slope is $0.9992$ ($R^{2}=1.0000$), confirming the Poisson relation $\sigma^{2}=\mu$ across two decades.
@@
Figure 6: C8: Quantum Harmonic Oscillator ($\sigma=\sqrt{\gamma\hbar\omega/m}$). (a) Simulated velocity time series $v(t)$ of the damped QHO (blue, in physical units) over $\sim$300 ps, with the pipeline-recovered local drift (red, $k$-NN with embedding dimension $m^{*}=5$). The drift smoothly tracks the oscillatory dynamics, capturing the reversal near $t\approx 150$ ps. (b) Probability density of locally recovered $\hat{\sigma}(x)$ values across the correlation manifold. The distribution is approximately Gaussian; the pipeline median $\hat{\sigma}=1.02\times 10^{5}$ (green line) agrees with the theoretical value $\sigma=\sqrt{\gamma\hbar\omega/m}=1.03\times 10^{5}$ (red dashed) to $0.35\%$. (c) Recovered $\hat{\sigma}$ versus theoretical $\sigma=\sqrt{\gamma\hbar\omega/m}$ on log–log axes across nine frequencies spanning one decade. Blue circles lie on the identity line (green, “Perfect agreement”); $R^{2}=0.9999$. The linear dependence $\sigma\propto\sqrt{\omega}$ confirms the fluctuation–dissipation relation at the quantum level. (d) Relative error (%) for 20 independent random seeds at the primary frequency. All errors lie below $1\%$; the mean error (blue dotted line) is $\sim$0.47%. The green dashed line marks zero error.
@@
Figure 7: C5: Van Kampen Scaling ($\sigma\propto\Omega^{-1/2}$). (a) Concentration time series $c=n/\Omega$ at system size $\Omega=500$ (blue) over 60 time units, with the 250-realisation ensemble median (green) and 95% CI band (green shading). Inset: probability densities of the true (blue) and forecast (green) concentration, showing close agreement. (b) Recovered diffusion profile $\hat{\sigma}(c)$ versus concentration (cyan scatter) with the theoretical state-dependent curve $\sigma(c)=\sqrt{(k_{b}/\Omega+k_{d}c)/\Omega}$ (red dashed). The scatter reflects local $k$-NN estimation; the theoretical curve passes through the centre of the cloud. (c) Recovered drift $\hat{\mu}(c)$ versus concentration (yellow scatter) with the theoretical mean-reversion $\mu(c)=k_{b}/\Omega-k_{d}c$ (red dashed). The equilibrium concentration $c_{\mathrm{eq}}=10$ is annotated; the drift changes sign at this point, confirming mean-reverting dynamics. (d) Median recovered $\hat{\sigma}$ versus system size $\Omega$ on log–log axes (cyan circles with 95% CI from 50 runs at each $\Omega$). The fitted slope $\alpha=-0.502\pm 0.001$ (solid pink) matches the Van Kampen prediction $\alpha=-0.500$ (dotted black); $R^{2}=1.0000$. (e) Distribution of the scaling exponent $\hat{\alpha}$ from 50 independent pipeline runs; median =0.501=-0.501 (green line), theoretical value 0.500-0.500 (red dashed), $p(\hat{\alpha}=-0.5)=0.458$, $N=50$. (f) Residuals $\Delta\log_{10}\hat{\sigma}$ from the fitted power law at each system size (brown circles). The RMS residual is $0.0011$ dex, confirming that the power-law relationship holds across two decades in $\Omega$.
@@
Figure 9: C9: QED Photon Field ($\sigma=c\sqrt{\gamma}$; $\hbar$ and $\omega$ cancel). (a) Dimensionless velocity $\tilde{v}$ of the photon mode versus dimensionless time $\omega t$ (blue), with the pipeline-recovered local drift (red, $k$-NN with $m^{*}=5$). The drift smoothly tracks the oscillatory dynamics over approximately four periods. (b) Probability density of locally recovered $\hat{\sigma}(x)$ in physical units ($\times 10^{14}$). The distribution is approximately Gaussian; the pipeline median (green line) of $\sim$$9.45\times 10^{13}$ agrees with the theoretical value $c\sqrt{\gamma}=9.48\times 10^{13}$ (red dashed) to $0.35\%$ (this single realisation; the mean error across 20 seeds is $0.50\%$, panel d). (c) The $\omega$-independence test. Recovered $\hat{\sigma}$ (blue circles) at seven frequencies spanning $\omega\in[10^{12},10^{13}]$ rad/s on a logarithmic frequency axis. The theoretical value $c\sqrt{\gamma}=9.480\times 10^{13}$ is marked (green horizontal line) with a $\pm 1\%$ band (green shading). All seven recovered values lie within the band and show no systematic trend with $\omega$, confirming the massless cancellation $\sigma=c\sqrt{\gamma}$. (d) Relative error (%) for 20 independent random seeds at the primary frequency $\omega=10^{13}$ rad/s (purple bars). The green dashed line marks zero error; the blue dotted line marks the mean error of $\sim$$0.50\%$. All but one realisation lies within $\pm 0.7\%$; the maximum excursion is $\sim$$1.0\%$ (realisation 8).

### 25 Mar 2026 v2

@@ 4.6 Relation to Nelson’s Stochastic Mechanics
Nelson [6] solved the forward problem for quantum mechanics: given $\sigma=\sqrt{\hbar/m}$, derive the Schrödinger equation. The present work solves the inverse problem across nine domains: given time series data, recover $\sigma$ without assuming the physics. The forward and inverse paths converge at C4, where the pipeline recovers Nelson’s $\sigma$ to $0.05\%$ (Table 2). Nelson’s insight, that physical law admits stochastic content [24], is here extended across the $\sigma$-continuum to gravitational degrees of freedom.
@@ 4.6.1 Distinguishing the stochastic and quantum formulations
The correspondence does not, however, extend to the level of solutions and physical predictions. Three structural properties of the Fokker–Planck evolution have no counterparts in the Wheeler–DeWitt equation.
First, the Fokker–Planck equation possesses a unique normalisable steady state. For any Fokker–Planck equation with non-degenerate diffusion and confining drift on a half-line or bounded domain, the zero-current condition $\partial_{g}[D\cdot P]=A\cdot P$ determines a unique normalisable distribution. The superspace Fokker–Planck equation (Appendix D, Eq. D.165), when coarse-grained to galactic-scale gravitational degrees of freedom via the Mori–Zwanzig projection (Appendix C, Proposition C.1), inherits this property: its steady state is uniquely determined by the drift and diffusion derived from A1–A4. The Wheeler–DeWitt equation $\hat{H}\Psi=0$ is a constraint equation: it has no time parameter, no spectral gap, and no mechanism for selecting a unique solution from its solution space.
Second, the Fokker–Planck evolution is dissipative. It drives any initial probability distribution toward the unique steady state on a timescale set by the spectral gap. This irreversible relaxation produces definite predictions: regardless of initial conditions, the distribution converges to the steady state. Quantum evolution preserves unitarity: pure states remain pure and the evolution is time-reversible: producing recurrences rather than relaxation.
Third, the Fokker–Planck steady state depends on the positivity of $P[g]\geq 0$ and the real-valued character of the diffusion coefficient. Under the Wick rotation, $P$ maps to a complex-valued wave functional $\Psi[g]$ that is generically oscillatory and not normalisable in the standard $L^{2}$ sense. The positive, real-valued, normalisable steady state of the Fokker–Planck equation has no natural image under this mapping.
@@
The framework therefore proposes a specific physical hypothesis: that the gravitational degrees of freedom evolve by diffusion on superspace, governed by the Fokker–Planck equation (Axiom A2), rather than by unitary quantum evolution governed by the Wheeler–DeWitt equation. The two formulations share a generator and a classical limit (general relativity, recovered as $L\gg\ell_{P}$ in both cases) but make distinct predictions for the statistical distribution of gravitational acceleration at galactic scales. The observational confrontation of these predictions with galactic kinematic data is the subject of subsequent work.
@@ 4.6.2 Relationship to conventional approaches to quantum gravity
The stochastic formulation is therefore not a reformulation of any existing approach to quantum gravity, but a distinct theoretical framework that shares their classical limit and connects to each through specific mathematical correspondences.
@@ 4.7 Limitations
All C1–C9 data are synthetic. The pipeline has not been applied to raw experimental data in this work. The synthetic data validates the methodology, where experimental application is the subject of subsequent publications. The blind protocol ensures that the pipeline cannot use domain knowledge, but the selection of the nine test domains is informed by the authors’ knowledge of physics; the pipeline’s performance on domains outside this selection is an open question.
The non-parametric $k$-NN Kramers–Moyal estimator carries a systematic positive bias of $O(k/N)$ in the recovered diffusion coefficient, ranging from $0.2\%$ (C8, C9) to $5.9\%$ (C3) across the nine domains. This bias is a mathematical property of local regression estimators, not a deficiency of the pipeline. It cancels in scaling exponents (demonstrated by C5) and is approximately uniform across state space for a given domain (demonstrated by the constant $\hat{\sigma}/\sigma_{\mathrm{true}}$ ratio across Lorentz factors in C7). The bias can be reduced by increasing $N$ or decreasing $k$, at the cost of increased variance. For the present study, the pipeline parameters ($d$, $k$, NN) were fixed per domain and not optimised to minimise the bias.
@@
The axioms A1–A4 are stated as postulates for deductive clarity, though each has independent justification. A1 adopts the standard canonical general relativity configuration space (Wheeler’s superspace [30]). A2 is a derived result (the Gravitational Diffusion Theorem, Appendix C) promoted to axiomatic status for deductive economy. A3 requires consistency with general relativity in the macroscopic limit. A4, the identification of probability as epistemic, is the sole genuinely free interpretive choice. Within A1–A4, the noise character, drift, covariance, fluctuation amplitude, and emergent time are all uniquely determined by theorem (Appendix D). Alternative axiomatisations may exist, but the physical content of the predictions follows from the derivation chain of Appendix C independently of the axiomatic packaging. The empirical question is therefore whether A1–A4 hold, not whether the predictions follow from them. If the axioms are incorrect, the predictions that follow from them need not be physically operative. This question is addressable by confronting the framework’s galactic-scale predictions with kinematic data.
To delimit scope explicitly: this paper does not claim to have derived quantum gravity, unified general relativity with quantum mechanics, or proved the axioms A1–A4. What is claimed is that if A1–A4 hold, then every element of the superspace SDE is uniquely determined as a non-parametric, first-principles consequence of the axioms. Whether A1–A4 are physically operative is an empirical question that can be addressed by observational data at galactic scales.
What the derivation chain does establish is that the gravitational diffusion coefficient is uniquely determined (the Gravitational Diffusion Theorem, Appendix C) by the same physical principles (the fluctuation–dissipation theorem, mass–energy equivalence, and the equivalence principle) that govern the nine validated domains, as a non-parametric, first-principles consequence of general relativity and quantum field theory.
@@ 5 Conclusion
Application of the resulting pipeline to nine physical domains (classical mechanics, statistical mechanics, nuclear physics, quantum mechanics, chemical kinetics, electromagnetism, relativistic quantum mechanics, quantum harmonic oscillator dynamics, and quantum electrodynamics) recovers the known governing equations of each domain blindly, with recovery errors ranging from $3\times 10^{-6}\%$ (deterministic domains) to approximately $6\%$ (absolute diffusion recovery at high $k/N$) and bootstrap confidence intervals quantifying the systematic $k$-NN estimator bias. For scaling exponents (C5), the multiplicative bias cancels and the bootstrap CI contains the theoretical value. Fundamental physical constants emerge in both the drift and the diffusion channels without prior specification. The recovered diffusion coefficients, viewed together, constitute an empirical pattern: the $\sigma$-continuum.
Three theorems derived from the relativistic and quantum field theory results yield three structural consequences. The derivation chain rests on the fluctuation–dissipation theorem (validated by C8 to $0.27\%$), mass–energy equivalence, and the equivalence principle, each among the most precisely tested results in physics. First, the Dambis–Dubins–Schwarz time-change theorem combined with the Klein–Gordon density produces the relativistic suppression $\sigma=\sigma_{0}/\gamma$, with physical constants emerging in both channels (Appendix B). Second, the massless cancellation and gravitational self-coupling together determine $\sigma=\ell_{P}/\sqrt{t_{P}}$ for gravitational degrees of freedom, with three independent consistency arguments converging on the dimensionless prefactor $\alpha=1$ (Appendix C). Third, a set of four canonical axioms (A1–A4) yields a framework within which physical time, the drift, the covariance operator, and the fluctuation amplitude are all uniquely fixed (Appendix D).
@@
is non-parametric within A1–A4: the noise character, drift, covariance, fluctuation amplitude, and emergent time are all uniquely determined by theorem (Appendix D). The empirical question is therefore whether A1–A4 hold, not whether the predictions follow from them.
The Superspace Diffusion Framework defined by axioms A1–A4 is falsifiable. The Fokker–Planck equation on superspace, coarse-grained via the Mori–Zwanzig projection to galactic-scale gravitational degrees of freedom, generates non-parametric, first-principles predictions for the statistical distribution of gravitational acceleration. These predictions are testable against galactic kinematic data including rotation curves, velocity dispersions, wide binaries, galaxy clusters, and gravitational lensing. If the axioms are incorrect, the predictions derived from them need not be physically operative: the observational confrontation therefore tests the axioms themselves, not merely their mathematical consequences.
@@ Acknowledgements
Figure 1: The $\sigma$-Continuum. Recovered diffusion coefficient $\sigma$ across all nine physical domains and the derived gravitational value, displayed on a logarithmic vertical axis ($\log_{10}\sigma$, schematic units). Each domain is represented by a distinct marker shape and colour, labelled with its recovered $\sigma$ expression. Deterministic domains (C1: classical mechanics, dark blue circle; C6: classical electromagnetism, light blue circle) are placed at $\log_{10}(\sigma)\approx-40$ for visualisation, labelled $\sigma=0$. Stochastic domains span intermediate scales: C2 (statistical mechanics, green square, $\sigma=\sqrt{2\gamma k_{B}T/m}$), C3 (radioactive decay, orange diamond, $\sigma^{2}=\mu$), C4 (quantum mechanics, purple pentagon, $\sigma=\sqrt{\hbar/m}$), C5 (chemical kinetics, red triangle, $\sigma\propto\Omega^{-1/2}$). The relativistic and quantum field theory domains occupy the upper region: C7 (Klein–Gordon, red–brown pentagon, $\sigma=\sqrt{\hbar/m}/\gamma$), C8 (quantum harmonic oscillator, brown diamond, $\sigma=\sqrt{\gamma\hbar\omega/m}$), C9 (QED photon, magenta circle, $\sigma=c\sqrt{\gamma}$). The dashed lilac horizontal line marks the quantum scale $\sigma\sim\sqrt{\hbar/m}$; the dashed gold horizontal line marks the Planck scale $\ell_{P}\approx 10^{-35}$ m. The gold star at the far right represents $\sigma=\ell_{P}$ for gravitational degrees of freedom, derived by the Gravitational Diffusion Theorem (§4; Appendix C). The continuum from $\sigma=0$ (C1, C6) through thermal, quantum, and relativistic scales (C2–C9) is an empirical pattern arising from nine independent blind recovery experiments; the gravitational endpoint $\sigma=\ell_{P}$ is derived from this pattern by the Gravitational Diffusion Theorem.
@@
Figure 5: C3: Radioactive Decay ($\sigma^{2}=\mu$, Fano =1=1). (a) Poisson count series at $\mu=3000$ (blue) over 500 measurements, with the 250-realisation ensemble median (green) and 95% CI band (green shading). Inset: probability densities of the data (blue), ensemble (green), and theoretical Poisson–Gaussian approximation (red). (b) Recovered $\hat{\sigma}^{2}$ (pipeline) versus mean count $\mu$ on log–log axes (magenta circles) across seven mean count levels from $\mu\approx 100$ to $\mu\approx 10\,000$. The dashed red line is the Poisson prediction $\sigma^{2}=\mu$. The log–log slope is $1.0002$ ($R^{2}=0.9997$). (c) Observed count distribution at $\mu=3000$ (blue histogram) overlaid with the fitted Gaussian ($\mu=3001$, $\sigma=55.1$; red curve). The measured Fano factor is $1.0109$, consistent with Poisson statistics ($F=1$). (d) Autocorrelation function of the count series at 50 lags. The red dashed lines mark the 95% significance band ($\pm 0.0196$). Three of 50 lags exceed the band ($\sim$2 expected by chance), confirming temporal independence, a necessary property of Poisson counting. (e) Distribution of Fano factor $F=\hat{\sigma}^{2}/\hat{\mu}$ from 50 independent simulations at $\mu=3000$; median =1.0000=1.0000 (green line), run-level 95% range $[0.9612,1.0295]$. This panel shows the raw variance/mean ratio; through the pipeline, $\hat{\sigma}/\sqrt{\mu}=1.059$ with bootstrap 95% CI [1.051, 1.062][1.051,\,1.062] (Table 2), reflecting the $O(k/N)$ estimator bias that cancels in the scaling-law test of panel (f). (f) Variance $\hat{\sigma}^{2}$ versus mean $\mu$ from 50-simulation medians at seven count levels (magenta circles with 95% CI); the log–log slope is $0.9992$ ($R^{2}=1.0000$), confirming the Poisson relation $\sigma^{2}=\mu$ across two decades.
@@
Figure 6: C8: Quantum Harmonic Oscillator ($\sigma=\sqrt{\gamma\hbar\omega/m}$). (a) Simulated velocity time series $v(t)$ of the damped QHO (blue, in physical units) over $\sim$300 ps, with the pipeline-recovered local drift (red, $k$-NN with embedding dimension $m^{*}=5$). The drift smoothly tracks the oscillatory dynamics, capturing the reversal near $t\approx 150$ ps. (b) Probability density of locally recovered $\hat{\sigma}(x)$ values across the correlation manifold. The distribution is approximately Gaussian; the pipeline median $\hat{\sigma}=1.02\times 10^{5}$ (green line) agrees with the theoretical value $\sigma=\sqrt{\gamma\hbar\omega/m}=1.03\times 10^{5}$ (red dashed) to $0.35\%$. With the embedding lag set to the autocorrelation zero-crossing ($\tau_{\mathrm{ACF}}$), 50 independent runs yield $\hat{\sigma}/\sigma_{\mathrm{true}}=0.997$ with bootstrap 95% CI [0.995, 1.000][0.995,\,1.000] (Table 2). (c) Recovered $\hat{\sigma}$ versus theoretical $\sigma=\sqrt{\gamma\hbar\omega/m}$ on log–log axes across nine frequencies spanning one decade. Blue circles lie on the identity line (green, “Perfect agreement”); $R^{2}=0.9999$. The linear dependence $\sigma\propto\sqrt{\omega}$ confirms the fluctuation–dissipation relation at the quantum level. (d) Relative error (%) for 20 independent random seeds at the primary frequency. All errors lie below $1\%$; the mean error (blue dotted line) is $\sim$0.47%. The green dashed line marks zero error. The 50-run bootstrap analysis (Table 2) gives $0.27\%$ error.
@@
Figure 7: C5: Van Kampen Scaling ($\sigma\propto\Omega^{-1/2}$). (a) Concentration time series $c=n/\Omega$ at system size $\Omega=500$ (blue) over 60 time units, with the 250-realisation ensemble median (green) and 95% CI band (green shading). Inset: probability densities of the true (blue) and forecast (green) concentration, showing close agreement. (b) Recovered diffusion profile $\hat{\sigma}(c)$ versus concentration (cyan scatter) with the theoretical state-dependent curve $\sigma(c)=\sqrt{(k_{b}/\Omega+k_{d}c)/\Omega}$ (red dashed). The scatter reflects local $k$-NN estimation; the theoretical curve passes through the centre of the cloud. (c) Recovered drift $\hat{\mu}(c)$ versus concentration (yellow scatter) with the theoretical mean-reversion $\mu(c)=k_{b}/\Omega-k_{d}c$ (red dashed). The equilibrium concentration $c_{\mathrm{eq}}=10$ is annotated; the drift changes sign at this point, confirming mean-reverting dynamics. (d) Median recovered $\hat{\sigma}$ versus system size $\Omega$ on log–log axes (cyan circles with 95% CI from 50 runs at each $\Omega$). The fitted slope $\alpha=-0.502\pm 0.001$ (solid pink) matches the Van Kampen prediction $\alpha=-0.500$ (dotted black); $R^{2}=1.0000$. (e) Distribution of the scaling exponent $\hat{\alpha}$ from 50 independent pipeline runs; median =0.502=-0.502 (green line), theoretical value 0.500-0.500 (red dashed). Bootstrap 95% CI on median: [0.503,0.500][-0.503,\,-0.500], containing the theoretical value (Table 2). (f) Residuals $\Delta\log_{10}\hat{\sigma}$ from the fitted power law at each system size (brown circles). The RMS residual is $0.0011$ dex, confirming that the power-law relationship holds across two decades in $\Omega$.
@@
Figure 9: C9: QED Photon Field ($\sigma=c\sqrt{\gamma}$; $\hbar$ and $\omega$ cancel). (a) Dimensionless velocity $\tilde{v}$ of the photon mode versus dimensionless time $\omega t$ (blue), with the pipeline-recovered local drift (red, $k$-NN with $m^{*}=5$). The drift smoothly tracks the oscillatory dynamics over approximately four periods. (b) Probability density of locally recovered $\hat{\sigma}(x)$ in physical units ($\times 10^{14}$). The distribution is approximately Gaussian; the pipeline median (green line) of $\sim$$9.45\times 10^{13}$ agrees with the theoretical value $c\sqrt{\gamma}=9.48\times 10^{13}$ (red dashed) to $0.35\%$ (this single realisation). With $\tau_{\mathrm{ACF}}$ embedding, 50 independent runs yield $\hat{\sigma}/\sigma_{\mathrm{true}}=0.998$ with bootstrap 95% CI [0.996, 0.999][0.996,\,0.999] (Table 2). (c) The $\omega$-independence test. Recovered $\hat{\sigma}$ (blue circles) at seven frequencies spanning $\omega\in[10^{12},10^{13}]$ rad/s on a logarithmic frequency axis. The theoretical value $c\sqrt{\gamma}=9.480\times 10^{13}$ is marked (green horizontal line) with a $\pm 1\%$ band (green shading). All seven recovered values lie within the band and show no systematic trend with $\omega$, confirming the massless cancellation $\sigma=c\sqrt{\gamma}$. (d) Relative error (%) for 20 independent random seeds at the primary frequency $\omega=10^{13}$ rad/s (purple bars). The green dashed line marks zero error; the blue dotted line marks the mean error of $\sim$$0.50\%$. The 50-run bootstrap analysis (Table 2) gives $0.24\%$ error.